OpenStudy (anonymous):

4.5x-3.8(greater than or equal to symbol)1.5x-2.3

OpenStudy (anonymous):

\[4.5x - 3.8 \ge 1.5x - 2.3\] Subtract 1.5x from both sides \[3x - 3.8 \ge -2.3\] add 3.8 to both sides \[3x \ge 1.5\] divide by 3 \[x \ge .5\]